a man can swim with a speed of 4km/hr in still water.how long does he take to cross a river 1km wide,if the river flows steadily at 3km/hr and he makes his strokes normal to the river current?how far in the river does he go,when he reaches the other bank?

Grade:12th pass

6 Answers

Mayank Kumar
7 years ago
let velocity of river be Vr = 3km/h
the time (t) taken to cross the river will be:- 
t= river width/speed of man   (time=dis/speed)
t=1/4 hour
 
Now dis=speed*time
distance moved in time t= Vr*t= 3km/h*1/4h=3/4 km
 
So distance travelled by man = 3/4km=750 meter
hope this helps.
Anshu Sharma
6 years ago
speed of man=4km/hwidth=1kmsince, time=distance/velocity time=1/4t=0.25hournow convet into minutes for easy callcuationt=0.25.60=15minutesvelocity of water=3km\hournow, due to motion of river water the man will reach the other bank of the river. distance=speed x timed=3x0.25=0.75kmdistance= 750m
Anjana
6 years ago
So basically all u need to use here is speed = distance /time formula.So the width of river = 3kmSpeed of man is 4km/hSo time taken by man to reach other end= width of river /speed of man.That is.... 1km/4km/h = 1/4 h= 15 minutes.Now the distance travelled by him is nothing but time taken by him * speed of river = 1/4h * 3km/h = 3/4km = 750 m
IItian
5 years ago
So try again you will definitely get the answer it's my promise jam a IITians
So believe in yourself use your brain it's an easy question guys let's do it
 
Manikho
4 years ago
Basically all the hints are given 
Swimmer speed =4.0km
River flowing speed =3.0km
Length of river =1km
So, time taken to cross river =dist./speed 
                                                  =1/4
Finally 
The dist. covered when the swimmer reached the other bank =speed ×time
                    =3×1/4
                     =0.75km or 750m
Simple as that.  Hope you find it useful 
                    =0.75km
      
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to your question below.
velocity of river be Vr = 3km/h
the time (t) taken to cross the river will be:- 
t = river width/speed of man   (time=dis/speed)
t = 1/4 hour
 
Now dis = speed*time
distance moved in time t = Vr*t = 3km/h*1/4h = 3/4 km
 
So distance travelled by man = 3/4km=750 meter
Hope it helps.
Thanks and regards,
Kushagra